OpenStudy (anonymous):

The coordinates below represent two linear equations. How many solutions does this system of equations have? Line 1 x y -6 3 3 6 Line 2 x y -3 1 3 3 A. 0 B. exactly 1 C. exactly 2 D. infinitely many

OpenStudy (anonymous):

@DanJS

OpenStudy (anonymous):

@sleepyjess

OpenStudy (anonymous):

I think its D

OpenStudy (anonymous):

Is that correct?

Nnesha (nnesha):

did you find the slope ??

Nnesha (nnesha):

if you find slope first then it would be easy to find answer

OpenStudy (anonymous):

k\[\frac{ -6 }{ 3 } and \frac{ -3 }{ 0 }\]

Nnesha (nnesha):

nope :( i guess you flipped that y values should at the top and x values should be at the bottom

OpenStudy (anonymous):

oh. k

OpenStudy (anonymous):

let me fix it

Nnesha (nnesha):

same like last one use same method to find slope you look tired but you can do it come on

Nnesha (nnesha):

okay sounds good :)

OpenStudy (anonymous):

\[-\frac{ 3 }{ 9 } and \frac{ 2 }{ }\]

OpenStudy (anonymous):

sorry there's a 6 under that 2

Nnesha (nnesha):

okay for first one numbers are good but negative sign be care full about that

Nnesha (nnesha):

 now solve this